# Designing a compact cavity-enhanced source of entangled photons

### v1

@@ Abstract
Quantum repeaters will require sources of entanglement allowing efficient coupling to quantum memories. Here, we address this challenge with a compact, narrowband source design. The entangled pairs are generated via SPDC in two perpendicularly oriented nonlinear crystals in a Fabry-Pérot cavity. We show that the case of highly non-degenerate wavelengths and type-II phase matching is the most promising candidate for a practical implementation. Using the parameters of an experiment we are currently working on, this design should allow generating entangled photons with a bandwidth of a few MHz.
@@
Large-scale quantum networks will use entanglement as a resource for future quantum applications Horodecki et al. (2009); Wehner et al. (2018). Establishing such networks is a challenge due to the losses in the optical links connecting the individual nodes. The distribution of entanglement has been demonstrated successfully over a few hundred kilometers in optical fiber networks Neumann et al. (2022); Zhuang et al. (2024). Using optical satellite links allows for the distribution over even larger distances, but at the cost of significant losses Yin et al. (2017); Sidhu et al. (2021); Kržič et al. (2023). In optical fibre networks, the range achievable for entanglement distribution can be increased by using quantum repeaters Briegel et al. (1998). These combine entanglement distillation and purification Bennett et al. (1996); Deutsch et al. (1996); Wootters (1998) with entanglement swapping Żukowski et al. (1993) to distribute high-fidelity entanglement over large distances. To avoid exponential loss, the entanglement needs to be stored efficiently until the intermediate steps have been completed successfully Bennett et al. (1996).
@@
Using narrowband sources of entanglement is beneficial for several purposes. It can increase the efficiency of storing entangled photons in quantum memories Wang et al. (2019); Heller et al. (2022) and it can reduce the background in the day-light operation of free-space links Aspelmeyer et al. (2003); Kržič et al. (2023). Moreover, using narrowband modes allows for frequency division multiplexing. This can increase the bandwidth of quantum channels and allows the simultaneous connection of multiple network clients Neumann et al. (2022). Generating narrowband photons can be achieved using cavity-enhanced spontaneous parametric downconversion (SPDC) Chuu et al. (2012); Lenhard et al. (2015). Other methods include spontaneous four-wave mixing in hot atomic vapor cells Zhu et al. (2017); Hsu et al. (2021) or using laser-cooled atoms Wang et al. (2022); Bruns et al. (2022). Recent works showed the generation of photon pairs compatible with the storage in atom-based quantum memories Tsai and Chen (2018); Prakash et al. (2019); He et al. (2023) and in solid-state quantum memories Liu et al. (2020); Onozawa et al. (2022). In some cases, the actual storage of single photons in quantum memories has already been achieved Mannami et al. (2021); Buser et al. (2022); Ito et al. (2023).
@@
Generating entanglement, however, requires making two biphoton processes indistinguishable. For example, by exciting indistinguishable processes in a single SPDC crystal Kwiat et al. (1995), or by creating a superposition of two-photon amplitudes originating in multiple crystals Kwiat et al. (1999); Ljunggren and Tengner (2005). Alternatively, the amplitudes can originate from multiple branches in an interferometer Fedrizzi et al. (2007). Recently, there has been significant progress in simultaneously achieving entanglement and narrow bandwidth of the generated photons. In Ref. Gao et al. (2024), the light from two cavity-enhanced SPDC crystals was used to generate entanglement with sub-GHz bandwidth in an interferometric setup. Refs. Niizeki et al. (2020); Ito et al. (2023) showed the generation of narrowband entanglement using a frequency comb. These photons then were frequency converted to a wavelength compatible with solid-state quantum memories. The authors of Ref. Arenskötter et al. (2023) used a single crystal in a bi-directionally pumped bow-tie cavity to generate entangled photons compatible with a ${}^{40}\mathrm{Ca}^{+}$ quantum memory.
@@
Here, we present the design of a compact, cavity-enhanced SPDC source of entangled photons without the need for complex interferometric or cavity setups. We will consider realistic experimental parameters to investigate the feasibility of achieving narrowband entanglement using a comparatively simple design. The source (Fig. 1) is based on a single cavity that contains two nonlinear crystals rotated by 90 degrees relative to each other, inspired by Ljunggren et al. (2006); Trojek and Weinfurter (2008); Chuu and Harris (2011); Chuu et al. (2012). The crystals can be of different lengths, $l_{1}$ and $l_{2}$. We pump the crystals with light at frequency $\omega_{p}$ and with the polarization chosen to balance the brightness of both SPDC processes.
@@
Figure 1: Schematic of two nonlinear crystals of lengths $l_{1}$ and $l_{2}$ in a Fabry Perót cavity. The green line indicates the pump beam, the blue and red lines indicate the signal and idler photons generated by SPDC. The optical axes of the crystals are mutually orthogonal.
@@
Depending on its polarization, the pump photon will create a pair of SPDC photons in the first or the second crystal. These photons are denoted as the signal and the idler with frequencies $\omega_{s}$ and $\omega_{i}$, respectively. Like the crystal axes, the polarization directions of the photons created in the first crystal will be perpendicular to those created in the second crystal.
@@
$\varphi$ is a relative phase and $a^{\dagger}_{n}(\omega)$ is a boson creation operator of mode $j$ at frequency $\omega$. $j$ can indicate the polarization or the crystal of origin. The polarization of the photons generated is determined by the PM condition. For type-II SPDC, the state will be a superposition of the Bell states $|\psi^{\pm}\rangle$, while it will be a superposition of Bell states $|\Phi^{\pm}\rangle$ for type-0 and type-I SPDC 111$|\psi^{\pm}\rangle=(1/\sqrt{2})\left(|HV\rangle\pm|VH\rangle\right)$ and $|\Phi^{\pm}\rangle=(1/\sqrt{2})\left(|HH\rangle\pm|VV\rangle\right)$. If the signal and idler photons originate from type-II SPDC and have degenerate frequencies, the amplitudes from different crystals will result in a product state $a^{\dagger}_{H}(\omega)a^{\dagger}_{V}(\omega)(1+e^{i\varphi})|\mathrm{vac}\rangle$. Using our source design with type-II PM to generate entanglement is therefore not possible for degenerate frequencies. For type-0 and type-I SPDC, one cannot spatially separate the photons if the frequencies are degenerate. For these reasons, we will focus on two specific cases: (1) the photons produced have near-degenerate wavelengths suitable for the low-loss distribution in telecommunication fibers, (2) only one photon is at such a telecommunication wavelength, and the second one has a wavelength suitable for efficiently coupling to a quantum memory.
Typically, the bandwidth of SPDC photons is much broader than the acceptance bandwidth of quantum memories Wang et al. (2019); Heller et al. (2022). The effective bandwidth can be reduced by coupling the process to an optical cavity Boyd and Ashkin (1966); Chuu and Harris (2011); Chuu et al. (2012); Tsai and Chen (2018). Only the portions of the photon spectra that are resonant with the narrow cavity modes will be enhanced. In comparison, the remainder of the spectra will be strongly suppressed. For each mode, denoted by an index $j$, we can define the mode number $m_{j}$, which is the phase shift that light in mode $j$ acquires during one cavity round trip. For example, in the case of a two-crystal type-II source, the four corresponding mode numbers are:
@@
Here, $m_{1}$ and $m_{2}$ represent the downconverted signal and idler mode numbers originating in the first crystal, while $m_{3}$ and $m_{4}$ represent the equivalent mode numbers for the second crystal. A single mode $j$ is resonant with the cavity if its frequency is an integer multiple of the free spectral range $\text{FSR}_{j}=\left|\partial_{\omega}m_{j}\right|^{-1}/2\pi$. Here, $\partial_{\omega}$ denotes the partial derivative with respect to $\omega$. To achieve narrowband entanglement between photons in these four modes, all of them have to be resonant with the cavity simultaneously. We denote this as a quadruple resonance.
Figure 2: (a) Illustration of the cluster spacings of double resonances and the joint cluster spacing for quadruple resonances. Red and black lines indicate idler and signal resonances, respectively. The upper segment shows the occurrence of double resonances for SPDC in the first crystal, the middle segment for SPDC in the second crystal. The bottom segment shows the occurrence of double resonances for both crystals, and the gray-shaded area highlights the occurrence of quadruple resonances, where the signal and idler modes from both SPDC processes are resonant simultaneously. (b) The fidelity of the state of a photon pair from a single crystal in well-defined frequency modes as a function of the cluster spacing normalized to the SPDC bandwidth.
Due to the frequency-dependent birefringence of the crystals, and due to their different lengths, the FSRs of the downconverted modes will be different. Simultaneous resonances of multiple modes will occur only at specific frequency spacings much greater than individual FSRs. This is known as cluster spacing. The cluster spacing $\Omega_{1}$ of a double resonance between modes $1$ and $2$ can be expressed in terms of their FSRs Chuu and Harris (2011); Chuu et al. (2012):
@@
One wants to have the joint cluster spacing small enough such that finding a quadruple resonance is achievable by tuning control parameters within a parameter range that is experimentally feasible. An additional challenge is finding the right ratio between the cluster spacing and the SPDC bandwidth. If the cluster spacing in an individual crystal is smaller than the corresponding SPDC bandwidth, multiple cavity modes at different frequencies will be enhanced. The output state of the cavity-enhanced SPDC in crystal $\alpha$ will then be a linear combination of these modes: $\sum_{j}A^{\alpha}_{j}a_{1}^{\dagger}(\omega_{s}+j\Omega_{\alpha})a_{2}^{\dagger}(\omega_{i}-j\Omega_{\alpha})|\mathrm{vac}\rangle$. Here, $A^{\alpha}_{j}=\mathrm{sinc}(l_{\alpha}\Delta k_{j}/2)$ are normalized amplitudes depending on the phase mismatch of the SPDC process $\Delta k_{j}$ (see Appendix A). $\alpha=1,2$ denotes crystal $1$ or $2$. Fig. 2(b) shows the calculated fidelity of the output state of crystal 1 for type-II PM with a product state $a_{H}^{\dagger}(\omega_{s})a_{V}^{\dagger}(\omega_{i})|\mathrm{vac}\rangle$ as a function of cluster spacing normalized to the SPDC bandwidth. The plot shows that a fidelity larger than 0.9 is achieved for
@@
This gives us a target bound for the source design if no additional spectral filtering is used.
In the following, we will assume realistic experimental parameters and analyze the feasibility of the source design for the cases of near-degenerate and highly non-degenerate frequencies using type-0 and type-II PM. For the SPDC crystals, we assume $5\%\,$MgO-doped periodically poled Lithium Niobate (PPLN) due to its high nonlinearity and transparency at $1550$ nm Gayer et al. (2008).
First, let us consider a near-degenerate source. We assume a pump wavelength of 775775775\,775nm and near-degenerate signal and idler modes at $21550\,\textrm{nm}\pm\Delta\lambda/$ 2. $\Delta\lambda$ corresponds to a small frequency shift $\Delta\omega=2\pi\times 80\,$MHz.

### v2

# thanks: These two authors contributed equallythanks: These two authors contributed equally

@@ Abstract
Entanglement will be the key resource of future large-scale quantum networks, enabling quantum communication and advanced quantum applications like distributed quantum sensing and distributed quantum computing. To this end, entanglement will have to be distributed over large distances and efficiently coupled to quantum devices at the network nodes. This requires the entangled photons to have wavelengths and bandwidths compatible with the quantum memories in quantum repeater nodes or quantum devices at client nodes. Here, we present a novel cavity-enhanced source design using two nonlinear crystals inside a single cavity. We provide detailed considerations balancing the complexity of the cavity design with the photon bandwidth and the entanglement quality.
@@
I Introduction
Large-scale quantum networks will use entanglement as a resource for future quantum applications [1, 2]. While the distribution of entanglement has been demonstrated over a few hundred kilometers in optical fiber networks [3, 4], the distribution over larger distances requires to either use quantum repeaters [5] or to use space optical links [6, 7, 8]. Quantum repeaters typically rely on the efficient storage of the distributed entanglement in quantum memories [9]. This requires the matching between the photon bandwidth and the memory linewidth [10]. Depending on the type of quantum memory used, this can impose limitations on the bandwidth of the entangled photons to allow their efficient storage. While some quantum memories allow for high bandwidths up to many GHz [11], others require a narrow bandwidth of a few MHz. For example, this is necessary to couple efficiently to the D2 transition of Rubidium [12], or to long-lived optomechanical memories [13].
To effectively use the distributed entanglement, it must be coupled to quantum devices at client nodes of the network. This can pose additional requirements on the entanglement. For example, the bandwidth has to be sub-MHz for optomechanical devices with mechanical frequencies in the MHz regime [13]. The bandwidth has to be even narrower if one wants to address a specific degree of motion of a trapped sub-micron particle. Such systems have, for example, become an attractive platform for realizing high-precision quantum sensors allowing for an unprecedented sensitivity in force measurements [14]. Coupling entangled photons to the center-of-mass motion of such particles will require photon bandwidths below $100$ kHz [15].
@@ I Introduction
Generating narrowband photons can be achieved using cavity-enhanced spontaneous parametric downconversion (SPDC) [16, 17, 18]. Other methods include spontaneous four-wave mixing in hot atomic vapor cells [19, 20] or using laser-cooled atoms [21, 22]. Recent works showed the generation of photon pairs compatible with the storage in atom-based [23, 24, 25] and solid-state [26, 27] quantum memories. The performance of quantum memories was tested on weak coherent pulses  [28, 29] and, in some cases, the actual storage of single photons has already been achieved [30, 31, 32].
@@
This shows there has already been significant progress in producing and storing narrowband biphotons. Generating entanglement between two photons additionally requires making two biphoton processes indistinguishable. For example, this can be achieved by exciting indistinguishable processes in a single SPDC crystal [33], or by creating a superposition of two-photon amplitudes originating in multiple crystals [34, 35]. Alternatively, the amplitudes can originate from multiple branches in an interferometer [36]. Early works showed the generation of entangled cavity-enhanced photons using two crystals inside one cavity [37] or by post-selecting an entangled state at the outputs of an interferometer [38, 39]. All these cases required significant spectral filtering to achieve narrowband entanglement. Rapid progress has been made in recent years in simultaneously achieving entanglement and a narrow bandwidth of the generated photons. In Ref. [40], the light from two cavity-enhanced SPDC crystals was used to generate entanglement with sub-GHz bandwidth in an interferometric setup. Refs. [41, 29] showed the generation of narrowband entanglement using a frequency comb. These photons then were frequency converted to a wavelength compatible with solid-state quantum memories. The authors of Ref. [42] used a single crystal in a bi-directionally pumped bow-tie cavity to generate entangled photons compatible with a ${}^{40}\mathrm{Ca}^{+}$ quantum memory.
In this paper, we present a novel design for a cavity-enhanced SPDC source. The entangled photons are generated inside a single cavity. While this will require a simultaneous resonance of four cavity modes, we will show that this is feasible with realistic experimental parameters. Even with a compact and simple cavity design, this will allow MHz linewidths with only moderate additional filtering. Significantly narrower bandwidths are possible if one allows for a more complex design. Additional dispersion compensation can be used to avoid the need for spectral filtering. A central goal of the paper is to investigate under which conditions realistic experimental parameters allow achieving high-quality entanglement. In Section II we present how entanglement can be generated in a two-crystal setup using SPDC. In Section III, we introduce the conditions for achieving simultaneous cavity resonances. We will apply these considerations in Section IV to two specific cases of phase matching. In Section V, we determine the parameters necessary for the SPDC crystals and the cavity in order to achieve a specific bandwidth. Finally, in Section VI we discuss which experimental parameters can be tuned to find and to lock the necessary cavity resonances.
II A two-Crystal Source of Entanglement
II.1 Spontaneous parametric downconversion
SPDC is a second-order nonlinear optical process in which a high-energy pump photon with frequency $\omega_{p}$ is annihilated and generates a pair of lower-energy photons with frequencies $\omega_{s}$ and $\omega_{i}$ [43]. These are denoted as the signal and the idler photons. Energy conservation results in $\hbar\omega_{p}=\hbar\omega_{s}+\hbar\omega_{i}$. The wavevectors of the photons have to fulfill phase-matching (PM) conditions for the process to occur with significant probability. In a periodically poled crystal, the PM conditions can be modified by the poling period of the crystal to achieve quasi-phasematching (QPM). Depending on the photon polarizations, the crystal type and its orientation, different PM configurations are possible. In type-I PM, the signal and idler photons have the same polarization, which is orthogonal to the polarization of the pump. In type-0 PM, all three photons have the same polarization. In type-II PM, the signal and idler photons have orthogonal polarizations.
The bandwidth of a single-pass SPDC process is determined by the phase mismatch $\Delta k=k_{s}+k_{i}-k_{p}-2\pi/\Lambda$. Here, $\Lambda$ is the period of the periodic poling of the SPDC crystal. The amplitude of the photons generated in a crystal of length $l$ is proportional to $(\sin\xi/\xi)^{2}$ with $\xi=\Delta kl/2$. The FWHM bandwidth $\Delta\nu_{\text{SPDC}}$ of the SPDC photons can then be determined from $\xi_{\text{HWHM}}\approx 1.39$. Expanding $\Delta k$ up to linear order in signal frequency and temperature yields:
Here, $T$ is the crystal temperature, which affects the refractive index and the poling of the nonlinear crystal [44, 45]. The SPDC bandwidth $2\Delta\omega_{s}$ can be determined from the solution of Eq. 1. In a standing-wave cavity, the length $l$ in that equation has to be replaced by twice the crystal length because the pump will pass each crystal twice in one round trip (see Fig. 1).
II.2 Generating Entanglement
@@
The source (Fig. 1) is based on a single cavity that contains two nonlinear crystals rotated by 90 degrees relative to each other, inspired by [46, 47, 48, 16]. In principle, this allows for a very compact cavity design if the crystal properties are chosen according to the considerations we will present here. The crystals can be of different lengths, $l_{1}$ and $l_{2}$ such that $l=l_{1}+l_{2}$. We pump the crystals with continuous-wave light at frequency $\omega_{p}$ and with the polarization chosen to balance the brightness of both SPDC processes.
@@
Figure 1: Schematic of nonlinear crystals of lengths $l_{1}$ and $l_{2}$ in a Fabry Perót cavity. The green line indicates the pump. Blue and red lines indicate the signal and idler photons generated by SPDC. The optical axes of the crystals are rotated $90^{\circ}$.
@@
Depending on its polarization, the pump will create a pair of SPDC photons in the first or the second crystal. Like the crystal axes, the polarization directions of the photons created in the first crystal will be perpendicular to those created in the second crystal.
@@
$\varphi$ is a relative phase and $a^{\dagger}_{j}(\omega)$ is a boson creation operator of mode $j$ at frequency $\omega$. $j$ can indicate the polarization or the crystal of origin. The polarization of the photons generated is determined by the PM condition. For type-II SPDC, the state will be a superposition of the Bell states $|\psi^{\pm}\rangle$, while for type-0 and type-I SPDC it will be a superposition of Bell states $|\Phi^{\pm}\rangle$ 111 $|\psi^{\pm}\rangle=(1/\sqrt{2})\left(|HV\rangle\pm|VH\rangle\right)$ and $|\Phi^{\pm}\rangle=(1/\sqrt{2})\left(|HH\rangle\pm|VV\rangle\right)$ are the four Bell states.. The feasibility of generating entanglement depends on the frequencies and polarizations of generated photons. If the signal and idler photons originate from type-II SPDC and have degenerate frequencies, the amplitudes from different crystals will result in a product state $a^{\dagger}_{H}(\omega)a^{\dagger}_{V}(\omega)(1+e^{i\varphi})|\mathrm{vac}\rangle$. Using our source design with type-II PM to generate entanglement is therefore not possible for degenerate frequencies. For type-0 and type-I SPDC, one cannot spatially separate the photons if the frequencies are degenerate. For these reasons, we will focus on two specific cases: (1) the photons produced have near-degenerate wavelengths suitable for the low-loss distribution in telecommunication fibers, (2) only one photon is at a wavelength appropriate for telecommunication fibers, while the second photon’s wavelength is chosen to allow for efficiently coupling to a specific type of quantum memory.
III Cavity-enhancement
III.1 Simultaneous resonances
Often, the bandwidth of SPDC photons is much broader than the acceptance bandwidth of quantum memories [50, 51]. The effective bandwidth can be reduced by coupling the process to an optical cavity [52, 53, 48, 16, 23]. Only the portions of the photon spectra that are resonant with the narrow cavity modes will be enhanced. In comparison, the remainder of the spectra will be strongly suppressed. We describe these cavity modes by defining a mode number $m_{j}$ for each mode $j$. It represents the phase shift that light in mode $j$ acquires during one cavity round trip. For a two-crystal type-II source, the four corresponding mode numbers are:
@@ III.1 Simultaneous resonances
Here, $m_{1}$ and $m_{2}$ represent the downconverted signal and idler mode numbers originating in the first crystal, while $m_{3}$ and $m_{4}$ represent the equivalent mode numbers for the second crystal. For a type-0 or type-I source, the corresponding mode numbers are:
$\begin{split}m_{1}(\omega_{s})&=\frac{\omega_{s}}{\pi c}\left(l_{1}n_{o}(\omega_{s})+l_{2}n_{e}(\omega_{s})+l_{\text{air}}\right),\\ m_{2}(\omega_{i})&=\frac{\omega_{i}}{\pi c}\left(l_{1}n_{o}(\omega_{i})+l_{2}n_{e}(\omega_{i})+l_{\text{air}}\right),\\ m_{3}(\omega_{s})&=\frac{\omega_{s}}{\pi c}\left(l_{1}n_{e}(\omega_{s})+l_{2}n_{o}(\omega_{s})+l_{\text{air}}\right),\\ m_{4}(\omega_{i})&=\frac{\omega_{i}}{\pi c}\left(l_{1}n_{e}(\omega_{i})+l_{2}n_{o}(\omega_{i})+l_{\text{air}}\right).\end{split}$ (4)
A single mode $j$ is resonant with the cavity if its frequency is an integer multiple of the free spectral range $\text{FSR}_{j}=\left|\partial_{\omega}m_{j}\right|^{-1}/2\pi$. Here, $\partial_{\omega}$ denotes the partial derivative with respect to $\omega$. To achieve narrowband entanglement between photons in these four modes, all of them have to be resonant with the cavity simultaneously. We denote this as a quadruple resonance.
Figure 2: (a) Illustration of the cluster spacings of double resonances and the joint cluster spacing for quadruple resonances. Red and black lines indicate idler and signal resonances, respectively. Top segment: occurrence of double resonances for SPDC in the first crystal. Middle segment: double resonances in the second crystal. Bottom segment: double resonances for both crystals, and the gray-shaded areas highlight quadruple resonances, where the signal and idler modes from both SPDC processes are resonant simultaneously. (b) The fidelity of the state of a photon pair from a single crystal in well-defined frequency modes as a function of the cluster spacing normalized to the SPDC bandwidth.
Due to the frequency-dependent birefringence of the crystals, and due to their different lengths, the FSRs of the individual downconverted modes will be different. Simultaneous resonances of multiple modes will occur only over specific frequency spacings much greater than individual FSRs. This is known as cluster spacing. The cluster spacing $\Omega_{1}$ of a double resonance between modes $1$ and $2$ can be expressed in terms of their FSRs [48, 16]:
@@
One wants the joint cluster spacing sufficiently small such that finding a quadruple resonance is achievable by tuning control parameters within a range that is experimentally feasible. An additional challenge is finding the right ratio between the cluster spacing and the SPDC bandwidth. If the cluster spacing in an individual crystal is smaller than the corresponding SPDC bandwidth, multiple cavity modes at different frequencies will be enhanced at the same time. The presence of multiple enhanced frequency modes will reduce the fidelity with the targeted two-photon state.
III.2 The state fidelity
In order to quantify the quality of the state produced in an individual crystal, we calculate its fidelity with the biphoton state we target: $a_{1}^{\dagger}(\omega_{s})a_{2}^{\dagger}(\omega_{i})|\mathrm{vac}\rangle$. The output state of the cavity-enhanced SPDC in crystal $\alpha$ will be a linear combination of all possible frequency combinations of signal and idler photons within the SPDC bandwidth: $\sum_{j}A^{\alpha}_{j}a_{1}^{\dagger}(\omega_{s}+j\Omega_{\alpha})a_{2}^{\dagger}(\omega_{i}-j\Omega_{\alpha})|\mathrm{vac}\rangle$. Here, $A^{\alpha}_{j}=\mathrm{sinc}(l_{\alpha}\Delta k_{j}/2)$ are normalized amplitudes depending on the phase mismatch of the SPDC process $\Delta k_{j,\alpha}$,
$\alpha=1,2$ denotes crystal $1$ or $2$. Fig. 2(b) shows the calculated fidelity of the output state of crystal 1 for type-II PM with a product state $a_{H}^{\dagger}(\omega_{s})a_{V}^{\dagger}(\omega_{i})|\mathrm{vac}\rangle$ as a function of cluster spacing normalized to the SPDC bandwidth. A fidelity above 0.9 is achieved for
@@ III.2 The state fidelity
A source fulfilling this bound for both crystals in the cavity will not require spectral filtering because only the processes at two frequencies with narrow bandwidths will be enhanced.
IV Comparing resonance spacings to SPDC bandwidth
In the following, we will assume realistic experimental parameters and analyze the feasibility of the source design for the cases of near-degenerate and highly non-degenerate frequencies using type-0 and type-II PM. The near-degenerate source is centered at $1550$ nm and is suitable for distribution of entangled pairs through optical fibers. For the highly non-degenerate source, we consider entangled photons at $1550$ nm and $780$ nm - the former for long-distance communication and for potentially coupling to optomechanical devices, the latter for coupling to Rubidium quantum memories. For the SPDC crystals, we assume $5\%\,$MgO-doped periodically poled Lithium Niobate (PPLN) due to its high nonlinearity and transparency at $1550$ nm [45, 44].
Our analysis starts by finding the PM condition for the chosen wavelengths for an experimentally feasible periodic poling and temperature of the crystal to find the lowest possible SPDC bandwidth. For these parameters, we calculate the cluster spacing and the SPDC bandwidth and evaluate them in relation to Eq. (7). Based on this, we fix the type of PM and calculate the joint cluster spacing. We find the ratio of crystal lengths that gives a reasonably low joint cluster spacing such that quadruple resonance can be achieved by tuning the experimental parameters. A similar analysis can be performed for entangled photons at other combinations of frequencies.
IV.1 Near-degenerate source
First, let us consider a near-degenerate source suitable for distributing entanglement over telecom fibres. We assume a pump wavelength of 775775775\,775nm and near-degenerate signal and idler modes at $21550\,\textrm{nm}\pm\Delta\lambda/$ 2. $\Delta\lambda$ corresponds to a small frequency shift $\Delta\omega=2\pi\times 80\,$MHz. The shift is chosen such that a typical AOM can shift the frequency of one of the photons, potentialy making the source degenerate at detection stage.